# Leveraging Content and Acoustic Representations for Speech Emotion Recognition

### v2

@@ IV-D2 Fine-tuning and evaluation
We use a batch size of with a learning rate of - and train the model for epochs. The hidden dimension of the two-layer classification head is set to be . The AdamW optimizer is used here as well. All the models, including the CARE and the baseline systems, utilize the same classification backend. Thus, the design allows fair comparison of the different representations. 444Code available at https://github.com/iiscleap/CARE.
@@ IV-E Performance of CARE
The results on the downstream datasets using representations from the proposed CARE model are shown in Table II. A comparative analysis with several recent studies is also presented. These baseline models are categorized into two groups based on the number of parameters used during inference: base models (parameter size M), and large models (M), which also include LLM based models. The following observations are made for each category:
@@ IV-E2 Large models
SONAR [29] is selected as the speech encoder in this category. For the six English-based datasets, the pre-trained English speech encoder555https://dl.fbaipublicfiles.com/SONAR/spenc.eng.pt is used, while the French666https://dl.fbaipublicfiles.com/SONAR/spenc.v3ap.fra.pt and German777https://dl.fbaipublicfiles.com/SONAR/spenc.v3ap.deu.pt speech encoders are utilized for the CaFE and EmoDB datasets, respectively. Similar to the CARE backend, the layer representations from the SONAR encoder are linearly combined and the classification head is trained on the downstream task. Although SONAR has nearly four times the parameter size of CARE, our proposed model outperforms SONAR across all datasets except CMU-MOSI.
@@ IV-E3 LLM based models
Two versions of SALMONN [30] (B and B)888https://huggingface.co/tsinghua-ee/SALMONN are considered as examples of LLM-based models.
@@
Although LLM-based models show strengths in in-domain emotion recognition datasets, their performance declines on out-of-domain tasks, indicating limited generalizability across diverse tasks and multilingual emotional speech. 2) CARE’s advantage over speech SSL models like WavLM, HuBERT, and data2vec is expected, given that these models are trained on non-emotional data (see Sec. V-F for a related experiment). 3) Notably, CARE outperforms the multilingual SONAR model on CaFE and EmoDB datasets although it is trained on English speech only. This showcases the generalizability of our pre-training technique to out-of-domain tasks in SER.
@@ IV-F Emotional Attribute Prediction
We use the concordance correlation coefficient (CCC) as the metric, which indicates the agreement between the predicted and the ground truth attribute scores for this dataset. Denoting the mean, variance of ground truth by , and predicted scores by , , the CCC is defined as
In Eq. 4, is the Pearson’s correlation coefficient between the ground truth and the predicted scores. For training the downstream model, the representations from CARE and other models are aggregated similar to the categorical datasets. This is followed by a two-layer regression head with as the hidden dimension and as the output dimension ( for each of the three attributes). The objective is to increase the CCC between the ground truth and the predicted values for each of the dimensions of valence, arousal and dominance. In other words, the loss used for each attribute is . The results for this dataset along with other baseline models are shown in Table.III. We note that for this task, the CARE embeddings achieve the best results in terms of the valence and dominance attributes, while the performance on arousal is marginally better for the SALMONN-B model.
@@ V-A Comparison with Baselines
The performance of CARE is further compared with two baselines based on the supervisory signals used during pre-training. The results are shown in Table IV.
Baseline : For each downstream dataset, PASE+ features are extracted and a classification network is trained to predict the emotion class of each utterance similar to CARE. The performance of CARE surpasses that of the acoustic supervisory signal by (relative) on average across the datasets. This improvement can be attributed to the larger parameter size of CARE compared to the PASE+ model.
Baseline : In this approach, transcripts are generated for each downstream dataset using the Whisper-large-v3 model [34] and subsequently processed by a pre-trained RoBERTa model. The internal representations from RoBERTa are linearly combined, followed by training a two-layer classification head. Notably, aside from the MELD and CMU-MOSI datasets, CARE outperforms this baseline for all the other datasets, despite the Whisper-RoBERTa system having approximately ten times more parameters. The Whisper-RoBERTa system’s performance on RAVDESS-Song, CaFE, and EmoDB is expected, as these datasets contain identical transcripts spoken or sung with varying emotions, making it difficult for a transcript-based system to distinguish between different emotional utterances.
@@ V-B Importance of the Two Encoders
Key takeaways: 1) The combination of the acoustic and common encoder representations outperforms the best performing SSL model (HuBERT) by (relative) on average for the datasets (Table II). Given the similar parameter count, this performance suggests an advantage of our pre-training approach. 2) For the three out-of-domain datasets, the acoustic-common combination fares better than its semantic counterpart. This is likely because these datasets contain same sentences with varied emotional expressions, making the acoustic features more effective for distinguishing between utterances. 3) Across all datasets, the combination of both encoders in CARE yields the highest performance, suggesting that while the individual encoder performances are comparable, they capture distinct characteristics of the speech signal.
@@ V-D Initialization of Acoustic and Common Encoders
Key takeaways: 1) The model’s performance decreases with data2vec initialization, likely due to data2vec’s lower baseline performance compared to HuBERT and WavLM (see Table II). An exception is the CMU-MOSI dataset, where this initialization improves over the WavLM initialized model by (relative). 2) The HuBERT-initialized model performs best on the RAVDESS-Song and Emo-DB datasets. Notably, HuBERT outperforms WavLM for these two out-of-domain datasets (Table II). 3) Initialization impacts the acoustic and common encoders less than the semantic encoder, as the latter requires alignment with text representations, making RoBERTa initialization essential.
@@ V-E Choice of Acoustic Targets
V-F Continued Pre-training of WavLM
@@
A key question in assessing the effectiveness of our proposed pre-training methodology is the influence of the MSP-PODCAST pre-training dataset. Since all self-supervised learning (SSL) models are trained on neutral data, their ability to accurately discern emotions from speech signals is typically limited. The emotion recognition performance of these SSL models when pre-trained on emotion datasets thus becomes crucial. To explore the impact of pre-training setup in the proposed CARE, we continued the pre-training of the publicly available WavLM-base model, using the MSP-PODCAST dataset. This was done following the WavLM pre-training procedure, with masked language modeling loss, for an additional steps (similar to the CARE). The results of this experiment are shown in Fig. 4, wherein the performance of the continually pre-trained WavLM model is denoted by WavLM-Cont. Pretrain. The performance of the combination of the common and acoustic encoders is also shown for comparison.
@@
Key takeaway: Continued pre-training improves WavLM-base performance on certain downstream tasks, like IEMOCAP. However, except for IEMOCAP, WavLM performs worse than CARE’s acoustic-common encoder combination. As all the models in Fig. 4 have the same size (M) during inference, this experiment highlights the benefits of our proposed distillation based pre-training.
V-G Visualization of Layers of CARE
@@
In order to interpret the pre-trained model representations learnt by the acoustic and semantic encoder of CARE, we probe the representations from each encoder. We use the English part of the Emotional Speech Dataset (ESD) [63] for this analysis. We form pairs of utterances, where both the utterances of a pair have the same emotional label in all cases and they are derived from two different speakers. A total of pairs are considered and the cosine similarities of the pooled representations (for transformer layers) are shown in Fig.5. The two panels indicate the settings where the speech content in the two utterances are the same (top-panel) or different (bottom-panel). Note that, the representations are obtained from the pre-trained model with unsupervised data, hence we do not expect the model to fully encode emotions. Following are the observations from this visualization:
1.
The semantic encoder in layer has slightly more similarity than the acoustic encoder for the set of pairs with same content (top-panel).
2.
When the spoken content is different, the acoustic encoder has higher similarity than the semantic encoder, indicating that the acoustic encoder is beneficial when the emotion information cannot be reliably predicted from the textual content of the audio.
@@ VI Summary
Limitations and future scope: The MSP-PODCAST dataset is used for pre-training CARE, which has only hours of emotional speech data. Training the model on a bigger corpus may improve the performance of the model further. Another limitation of this work, is the relatively lower performance on some in-domain speech datasets compared to LLM-based models, like the CMU-MOSI dataset. In future, we plan to extend the CARE approach to multi-modal speech-text emotion recognition tasks.

### v3

@@ IV-D2 Fine-tuning and evaluation
We use a batch size of with a learning rate of - and train the model for epochs. The hidden dimension of the two-layer classification head is set to be . The AdamW optimizer is used here as well. All the models, including the CARE and the baseline systems, utilize the same classification backend. Thus, the design allows fair comparison of the different representations. 111Code available at https://github.com/iiscleap/CARE.
@@ IV-E Performance of CARE
The results on the downstream datasets using representations from the proposed CARE model are shown in Table II. These baseline models are categorized into two groups based on the number of parameters used during inference: base models (parameter size M), and large models (M), which also include LLM based models. The following observations are made for each category:
@@ IV-E2 Large models
SONAR [31] is selected as the speech encoder in this category. For the six English-based datasets, the pre-trained English speech encoder222https://dl.fbaipublicfiles.com/SONAR/spenc.eng.pt is used, while the French and German speech encoders are utilized for the CaFE and EmoDB datasets, respectively. Similar to the CARE backend, the layer representations from the SONAR encoder are linearly combined and the classification head is trained on the downstream task. Although SONAR has nearly four times the parameter size of CARE, our proposed model outperforms SONAR across all datasets except CMU-MOSI.
@@ IV-E3 LLM based models
Two versions of SALMONN [32] (B and B)333https://huggingface.co/tsinghua-ee/SALMONN are considered as examples of LLM-based models.
@@
Although LLM-based models show strengths in in-domain emotion recognition datasets, their performance declines on out-of-domain tasks, indicating limited generalizability across diverse tasks and multilingual emotional speech. 2) CARE’s advantage over speech SSL models like WavLM, HuBERT, and data2vec is expected, given that these models are trained on non-emotional data (see Sec. V-G for a related experiment). 3) Notably, CARE outperforms the multilingual SONAR model on CaFE and EmoDB datasets although it is trained on English speech only. This showcases the generalizability of our pre-training technique to out-of-domain tasks in SER.
@@ IV-F Emotional Attribute Prediction
We use the concordance correlation coefficient (CCC) as the metric. Denoting the mean, variance of ground truth by , and predicted scores by , , the CCC is defined as
In Eq. 5, is the Pearson’s correlation coefficient between the ground truth and the predicted scores. For training the downstream model, the representations from CARE and other models are aggregated similar to the categorical datasets. This is followed by a two-layer regression head with as the hidden dimension and as the output dimension ( for each of the three attributes). The objective is to increase the CCC between the ground truth and the predicted values for each of the dimensions of valence, arousal and dominance. The results for this dataset along with other baseline models are shown in Table.III. We note that for this task, the CARE embeddings achieve the best results in terms of the valence and dominance attributes, while the performance on arousal is marginally better for the SALMONN-B model.
@@ V Discussion
| Datasets | PASE+ [23] | Whisper [36] |
CARE | ||||||||
| Params:M | Params:M | Params:B | Params:B | Params:M | Params:M | Params:M | |||||||||
| IEMOCAP-4 [45] | |||||||||||||||
| IEMOCAP-6 [45] | |||||||||||||||
| MELD [46] | |||||||||||||||
| CMU-MOSI [47] | |||||||||||||||
| DAIC-WOZ [48, 63] | |||||||||||||||
| RAVDESS-Song [49] | |||||||||||||||
| CaFE [50] | |||||||||||||||
| EmoDB [51] | |||||||||||||||
| Avg. |
@@ V-A Comparison with Baselines
Four baseline systems (Table IV) are considered:-
PASE+: For each downstream dataset, PASE+ features are extracted and a classification network is trained to predict the emotion class of each utterance similar to CARE. The total number of parameters used during inference is M.
Whisper: For each downstream dataset, the representations from the encoder layers of the Whisper-large-v3 model [36] are linearly combined with learnable weights. A two-layer classification head is trained on top of these representations for the task of emotion recognition. The total number of parameters used during inference is M.
Whisper+RoBERTa: The transcripts are generated using the Whisper-large-v3 model and subsequently processed by a pre-trained RoBERTa model. The internal representations from RoBERTa are linearly combined by learnable weights, followed by training a two-layer classification head. This has a total of B parameters in use during inference.
Teacher-fusion: The PASE+ and Whisper+RoBERTa representations are concatenated and a two-layer classification head is trained for each downstream dataset. This baseline also has a total of B parameters during inference.
Key takeaways: 1) The performance of CARE surpasses that of the acoustic supervisory signal by (relative) on average across the datasets. This improvement can be attributed to the larger parameter size of CARE compared to the PASE+ model. 2) CARE is seen to outperform Whisper and Whisper+RoBERTa systems by and in relative terms. This indicates that, although the Whisper-based baselines are much larger in size, the combination of the acoustic and semantic information in CARE results in effective emotion recognition. 3) On MELD and CMU-MOSI, CARE is outperformed by the Whisper+RoBERTa baseline. For these datasets, text-based models are known to significantly outperform speech-only systems [19, 64].
In the Whisper+RoBERTa setup, the RoBERTa model is fine-tuned on transcripts generated by Whisper-large-v3 (B sized model). In contrast, CARE is a smaller model (M), and does not use directly use the ASR transcripts during inference. To further elucidate the fairness in model-size,
we replace Whisper-large-v3 with a Whisper-base model for the ASR, followed by the RoBERTa modeling. Then, the performance drops from to on MELD and from to on CMU-MOSI. This underscores the importance of accurate transcriptions and large model capacity in settings where the textual information is emotion rich. 4) While the teacher-fusion baseline is competitive for a number of datasets involving English speech, CARE outperforms this baseline on average by absolute. This also motivates why CARE was pre-trained using knowledge distillation as it outperforms the fusion baseline with only of the parameters.
@@ V-B Importance of the Two Encoders
Key takeaways: 1) The combination of the acoustic and common encoder representations outperforms the best performing SSL model (HuBERT) by (relative) on average for the datasets (Table II). Given the similar parameter count, this performance suggests an advantage of our pre-training approach. 2) For the three out-of-domain datasets, the acoustic-common combination fares better than its semantic counterpart.
3) Across all datasets, the combination of both encoders in CARE yields the highest performance, suggesting that while the individual encoder performances are comparable, they capture distinct characteristics of the speech signal.
@@ V-D Initialization of Acoustic and Common Encoders
Key takeaways: 1) The model’s performance decreases with data2vec initialization, likely due to data2vec’s lower baseline performance compared to HuBERT and WavLM (see Table II). An exception is the CMU-MOSI dataset, where this initialization improves over the WavLM initialized model by (relative). 2) The HuBERT-initialized model performs best on the RAVDESS-Song and Emo-DB datasets. Notably, HuBERT outperforms WavLM for these two out-of-domain datasets (Table II). 3) Initialization impacts the acoustic and common encoders less than the semantic encoder, as the latter requires alignment with text representations.
@@ V-E Choice of Acoustic Targets
V-F Choice of Semantic Targets
We conduct an experiment where the Whisper encoder representations serve as supervisory signals for the semantic encoder. We explore two variants of this: 1) We pool the Whisper representations to serve as semantic targets while pre-training. This model is called CARE (Whisper-pool). 2) We pre-train a model with the frame-level representations of Whisper as the targets. We call this model CARE (Whisper-frame). 3) We also use the frame level alignments between speech and the RoBERTa tokens and use the frame-level RoBERTa representations as the semantic targets. We call this model CARE (RoBERTa-frame). The comparative performances of the different variants along with the proposed model-CARE (RoBERTa-pool) are shown in Fig. 4.
Key takeaways: 1) The performance of CARE (RoBERTa-pool) is seen to be superior to both variants trained with Whisper encoded representations.
2) The performance of the systems when the semantic encoder is trained with the pooled targets is observed to be better than those trained with frame-level representations.
V-G Continued Pre-training of WavLM
@@
Since all self-supervised learning (SSL) models are trained on neutral data, their ability to accurately discern emotions from speech signals is typically limited. The emotion recognition performance of these SSL models when pre-trained on emotion datasets thus becomes crucial. To explore the impact of pre-training setup in the proposed CARE, we continued the pre-training of the publicly available WavLM-base model, using the MSP-PODCAST dataset. This was done following the WavLM pre-training procedure, with masked language modeling loss, for an additional steps (similar to the CARE). The results of this experiment are shown in Fig. 5, wherein the performance of the continually pre-trained WavLM model is denoted by WavLM-Cont. Pretrain. The performance of the combination of the common and acoustic encoders is also shown for comparison.
@@
Key takeaway: Continued pre-training improves WavLM-base performance on certain downstream tasks, like IEMOCAP. However, except for IEMOCAP, WavLM performs worse than CARE’s acoustic-common encoder combination. As all the models in Fig. 5 have the same size (M) during inference, this experiment highlights the benefits of our proposed distillation-based pre-training.
V-H Multimodal Emotion Recognition
To assess CARE’s utility in the multimodal speech-text setting, we design a model using speech (WavLM-base or CARE), and text (RoBERTa) fusion. After combining the layer representations in SUPERB style, we concatenate the representations and train a classification head. We experiment on IEMOCAP-4 and IEMOCAP-6 and observe that the weighted F1-score for CARE+RoBERTa improves from to for IEMOCAP-4 and from to for IEMOCAP-6 over WavLM-base+RoBERTa system. In addition to the uni-modal improvements reported in Table IV, these results highlight that multi-modal speech-text emotion recognition systems can also benefit from the enhanced representations provided by CARE.
V-I Visualization of Layers of CARE
@@
In order to interpret the pre-trained model representations learnt by the acoustic and semantic encoder of CARE, we probe the representations from each encoder. We use the English part of the Emotional Speech Dataset (ESD) [66] for this analysis. We form pairs of utterances, where both the utterances of a pair have the same emotional label in all cases and they are derived from two different speakers. A total of pairs are considered and the cosine similarities of the pooled representations (for transformer layer ) are shown in Fig. 6. The figure on the left indicates the setting where the speech content in the two utterances is the same whereas the plot on the right indicates different speech content. We note that when the spoken content is different, the acoustic encoder has higher similarity than the semantic encoder, indicating that the acoustic encoder is beneficial when the emotion information cannot be reliably predicted from the textual content of the audio.
@@ VI Summary
Limitations and future scope: The MSP-PODCAST dataset is used for pre-training CARE, which has only hours of emotional speech data. Another limitation of this work, is the relatively lower performance on some in-domain speech datasets compared to LLM-based models, like the CMU-MOSI. In future, we plan to extend the CARE approach to multi-modal speech-text emotion recognition tasks.